OpenStudy (anonymous):

IM DYING this is the last question someone please help !! (Cos theta - cos theta )^2 + (cos theta + cos theta)^2

OpenStudy (anonymous):

@Hero !!! Please I need your intelligence!!

OpenStudy (anonymous):

@mathmale !!

OpenStudy (anonymous):

\[\cos^2(θ)+\cos(θ)\]

OpenStudy (jdoe0001):

hmmm

OpenStudy (campbell_st):

well expand them they are both perfect squares (a + b)^2 = a^2 + 2ab + b^2 (a - b)^2 = a^2 - 2ab + b^2 that's all that is needed, then collect like terms

OpenStudy (jdoe0001):

\(\bf (a-a)^2+(a+a)^2\implies ?\)

OpenStudy (anonymous):

Can you explain please? My teacher gets irritated when I don't understand assignments

OpenStudy (anonymous):

The first two would cancel right ?

OpenStudy (jdoe0001):

yes

OpenStudy (campbell_st):

but the question doens't make a lot of sense since the 1st brackets is basically (a - a)^2 = 0 and the 2nd is (a + a)^2 = 4a

OpenStudy (jdoe0001):

well 4\(a^2\) but essentially that's all there is

OpenStudy (anonymous):

Would it be 4cos^2 theta ?????

OpenStudy (larseighner):

That is what it looks like, by I am not sure you stated your problem correctly.